OpenStudy (anonymous):

The given lengths are two sides of a right triangle. All three side lengths of the triangle are integers, and together they form a Pythagorean triple. Find the length of the third side, then indicate whether it is a leg or a hypotenuse. 28, 96 A. 37, leg B. 54, leg C. 92, leg D. 100, hypotenuse

OpenStudy (anonymous):

If the answer is an integer it means you get no decimal answers. So try plugging in these numbers into the Pythagorean Theorem and solve for the missing value. If you get a decimal try another combination.

OpenStudy (anonymous):

i tried

OpenStudy (anonymous):

Okay try, a and b are the legs. c is the hypotenuse. Let 28=a and 96=b. Plug them into the Pythagorean Theorem: \[a^2+b^2=c^2\] Solve for C

OpenStudy (anonymous):

ohh ok thanks mate

OpenStudy (anonymous):

No problem. Let me know if you need help.

OpenStudy (anonymous):

so its 784 for the first one and 9216 for the second one then do u add the together?

OpenStudy (anonymous):

them*

OpenStudy (anonymous):

Yes add them together. Then you will have to square root them since you have to get rid of the 2 on: \[c^2\]

OpenStudy (anonymous):

is it c? 92 leg

OpenStudy (anonymous):

No, go back to where you add them. 784+9216=10,000 \[10,000=c^2\]

OpenStudy (anonymous):

Then square root both sides. \[\sqrt{10,000}=\sqrt{c^2}\]

OpenStudy (anonymous):

k

OpenStudy (anonymous):

so the sq root of 10,000 is 100

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

then you multiply?

OpenStudy (anonymous):

Nope that will be your final answer. c=100 c is the hypotenuse of a triangle. So your answer is D.

OpenStudy (anonymous):

oh wow thank you so much

OpenStudy (anonymous):

Your very welcome. Feel free to ask for any more help if you need it. I will be getting off in 20 minutes but there's plenty of other members to help out.

OpenStudy (anonymous):

are you a teacher?

OpenStudy (anonymous):

Nope just a student on break. :)